OpenStudy (anonymous):

PLEASE HELP ME!!!! (-2 + 9i) + (9 + i)

OpenStudy (e.mccormick):

Associative Law of Addition. Order does not matter when there are only things being added. So reorder it and add.

OpenStudy (e.mccormick):

1+(2+3)=(1+2)+3=1+3+2=2+1+3 and so on.

OpenStudy (anonymous):

i dont understand